OpenStudy (anonymous):

use reference angle to find the exact value of the expression. tan 7pie/6

OpenStudy (anonymous):

\[\tan \frac{ 7\pi }{6 }=\tan \left( \pi+\frac{ \pi }{ 6 } \right)=\tan \frac{ \pi }{ 6 }=\frac{ \frac{ 1 }{2 } }{\frac{ \sqrt{3} }{2 } }=\frac{ 1 }{ \sqrt{3} }\]

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

yw